OpenStudy (anonymous):

Solve dy/dx=xy with initial condition y(0)=1.

OpenStudy (anonymous):

it is a separable ordinary differential equation. first step is to separate the y terms and x terms. \[\frac{ dy }{ y } = x dx\] next step is integrating both sides: \[\int\limits_{}^{} \frac{ dy }{ y } = \int\limits_{}^{}x dx\] \[\ln|y| + C = \frac{ x^2 }{ 2 } + C\] Both C's are arbitrary so you can add them and put them both on one side. \[\ln|y| = \frac{ x^2 }{ 2 } + C\] next step is to solve for y, wanna do it?

OpenStudy (anonymous):

y=e^(x^2/2), right?

OpenStudy (anonymous):

yes, but actually: \[y = e^{(\frac{ x^2 }{ 2 } + C)} = e^{\frac{ x^2 }{ 2 }}e^{C} \] Since e^C can be any real positive number, we can say that e^C = C (an arbitrary constant) , giving: \[y = Ce^{\frac{ x^2 }{ 2 } }\] this is the general solution for the form of differential equation that was solved "dy/dx = xy" the value of C depends on the initial condition. which in this case is y(0) = 1 plugging that in the general solution: \[y(0) = 1 = Ce^{0} = C\] C = 1 for this initial condition. Therefore the final solution is: \[y(x) = e^{\frac{ x^2 }{ 2 }}\]

OpenStudy (anonymous):

So the answer is y=e^(x^2/2).

OpenStudy (anonymous):

yes. sorry for delay lol

OpenStudy (anonymous):

$$\frac{dy}{dx}=xy\\\frac1ydy=x\,dx\\\int\frac1ydy=\int x\,dx\\\log y=\frac12x^2+C\\y=Ce^{\frac12x^2}$$

OpenStudy (anonymous):

$$y(0)=1\\Ce^0=1\\C=1$$so we know our particular solution is \(y=e^{\frac12 x^2}\)

OpenStudy (anonymous):

Thank you so much.